which type of line matches these equations x+y=10 and x-y=6
Answer by DrBeeee(684)

Im not sure what is is meant by type of line, but here goes.
Rewrite the two lines in slope/intercept form
(1) y = mx + b
You get two lines
(2) y = -x + 10 and
(3) y = x - 6
Note that the slope of (2), the coefficient of x, is -1 and the slope of (3) is +1. This gives us
(4) m1*m2 = -1 which proves that the two lines are perpendicular to each other.
If you solve (2) and (3) simultaneously you find that the lines cross at (8,2)